# An Information-Theoretic Approach to Identifying Formulaic Clusters in Textual Data

### v1

@@ Abstract
Texts, whether literary or historical, exhibit structural and stylistic patterns shaped by their purpose, authorship, and cultural context. Formulaic texts—characterized by repetition and constrained expression—tend to have lower variability in self-information compared to more dynamic compositions. Identifying such patterns in historical documents, particularly multi-author texts like the Hebrew Bible provides insights into their origins, purpose, and transmission. This study aims to identify formulaic clusters—sections exhibiting systematic repetition and structural constraints—by analyzing recurring phrases, syntactic structures, and stylistic markers. However, distinguishing formulaic from non-formulaic elements in an unsupervised manner presents a computational challenge, especially in high-dimensional and sample-poor data sets where patterns must be inferred without predefined labels.
To address this, we develop an information-theoretic algorithm leveraging weighted self-information distributions to detect structured patterns in text. Our pproach directly models variations in sample-wise self-information to identify formulaicity. By extending classical discrete self-information measures with a continuous formulation based on differential self-information in multivariate Gaussian distributions, our method remains applicable across different types of textual representations, including neural embeddings under Gaussian priors.
@@ 1 Introduction
An intriguing challenge in textual analysis is distinguishing formulaic layers—those characterized by recurring patterns, motifs, or stylistic conventions—from non-formulaic ones. In this context, a formulaic element is defined as a word or a sequence thereof, syntactic or grammatical structures, or thematic patterns that recur systematically within a corpus, either due to conventionalized linguistic usage, genre-specific constraints, or cultural transmission mechanisms. Formulaic expressions can function as structural scaffolds within texts, shaping their composition and reflecting underlying conventions of specific literary traditions, genres, or modes of authorship (e.g., Read and Nation, 2008; Paquot and Granger, 2012; Wood, 2019). Beyond their structural significance, analyzing formulaic patterns provides insights into the dynamic interplay between creativity and constraint in language, illuminating the socio-cultural and cognitive forces that have influenced textual transmission and stylistic evolution (e.g., Magoun Jr, 1953; Jensen, 1980).
@@
The importance of formulaicity is particularly pronounced in historical texts, where composition layers often reflect complex authorship and redaction processes. Many historical documents, created and transmitted over centuries, are products of collective memory and institutional influence. As such, these texts often contain formulaic structures that encode cultural norms, narrative conventions, and theological or political ideologies (e.g., Knohl, 2007; Polak, 2017). Understanding these layers can reveal how texts were constructed, transmitted, and adapted over time (e.g., Gitay, 1980; Polak, 2006; Coleman, 2019).
@@
Among historical texts, the Hebrew Bible provides a compelling case study. As a composite text shaped by centuries of authorship, redaction, and transmission, the Hebrew Bible embodies a rich interplay of formulaic and non-formulaic elements. Its stylistic diversity reflects contributions from different schools of authorship, genres, and literary motifs, ranging from tightly structured legal codes to more fluid poetic and narrative traditions (Wärters, 1976; Shectman and Baden, 2009; Stipp, 2017). Detection and analysis of formulaic patterns within the Hebrew Bible offer a pathway to better understanding its compositional history, shedding light on how diverse voices and traditions coalesced into the canonical form.
@@
Understanding the distinction between formulaic and non-formulaic layers requires rigorous methodologies to quantify stylistic and structural patterns across literary feature sets. One promising metric for addressing these challenges is entropy—a concept from information theory that measures the unpredictability or randomness of a dataset. Initially introduced by Rudolf Clausius in 1865 as a measure of energy dispersal in a system, and subsequently formalized to its statistically by Shannon (1948), entropy has found applications in diverse fields, including thermodynamics, cryptography, and data compression. In the context of textual analysis, entropy provides a framework for quantifying the variability and regularity inherent in linguistic structures. Formally, given a discrete probability distribution $P=\{p_{1},p_{2},\dots,p_{n}\}$ over $n$ possible elements, Shannon entropy is defined as:
@@
where $p_{i}$ represents the probability of the $i$th element. Higher entropy values indicate greater variability, while lower entropy values correspond to constrained, repetitive structures—making entropy a powerful tool for distinguishing formulaic clusters from more unpredictable distributions.
The relevance of entropy to textual analysis lies in its ability to capture the degree of predictability of literary or linguistic features within a sequence, such as a sentence, paragraph, or chapter. Texts with high formulaicity—such as legal codes or liturgical compositions—are likely to exhibit low entropy, such that text units within them resemble by, for example, vocabulary or grammatical structure. Conversely, more creative or extemporaneous texts often exhibit higher entropy, reflecting their greater unpredictability. information-based metrics, therefore, serve as a natural tool for distinguishing between formulaic and non-formulaic elements within texts (e.g., Church and Hanks, 1990).
@@
where $w_{i}$ represents the $i$th word in the sequence, and $P(w_{i}\mid w_{i-1},w_{i-2},\dots)$ denotes the conditional probability of $w_{i}$ given its preceding context. Perplexity has proven instrumental in analyzing textual corpora over various linguistic angles (e.g., Klakow and Peters, 2002; Huang and Hansen, 2007; Gamallo et al., 2017). In the context of formulaic analysis, perplexity can reveal how tightly a text adheres to specific linguistic conventions (e.g., Kurzynski, 2023).
@@
For instance, Gaussian Mixture Models (GMMs) (Vlassis and Likas, 2002) assume that the dataset can be represented as a weighted sum of Gaussian components, enabling efficient clustering through iterative optimization of parameters such as means, covariance matrices, and mixing weights. Similarly, Cross-Entropy Clustering (CEC) (Tabor and Spurek, 2014) explicitly incorporates entropy as a criterion to partition data, refining clusters based on their statistical coherence. These methods, while effective in many contexts, rely on assumptions about the underlying data distribution, which may not always align with the structural and stylistic non-Gaussian variability inherent in textual corpora.
Recent advances in computational models, particularly neural embeddings, have transformed the study of texts by representing textual units as vectors in a high-dimensional feature space (e.g., Devlin et al., 2018). Many of these models, including variational autoencoders (VAEs), assert that latent features follow a Gaussian distribution in a continuous space (Dilokthanakul et al., 2016; Yang et al., 2019; Li et al., 2020). VAEs excel in capturing the underlying structure of data by encoding inputs into compact latent representations governed by a Gaussian prior (Casale et al., 2018). This enables smooth interpolations and the generation of new data points resembling the original distribution.
@@
While these approaches provide useful frameworks for clustering and representation learning, their reliance on Gaussian assumptions and metrics that require estimation of covariance presents notable challenges. In textual data, where features often exhibit categorical, skewed, or multimodal distributions, the Gaussian prior may not accurately capture structural variability. Moreover, many of these methods are highly sensitive to the sample-to-dimension ratio (Altonji and Segal, 1996; Yao et al., 2015). When the sample size is small relative to the feature space, covariance estimation becomes unstable, leading to poorly conditioned matrices or singularities (e.g., Ashurbekova et al., 2021). Regularization techniques, such as adding small positive constants to the diagonal of covariance matrices, can partially mitigate these issues but introduce biases that require careful tuning (Bickel and Levina, 2008). This bias arises because regularization alters the estimated eigenvalues of the covariance matrix, effectively shrinking them toward a predetermined value, which can distort the true structure of the data and influence clustering or classification outcomes.
@@
These challenges are particularly pronounced in textual analysis, where statistical significance often necessitates dividing texts into relatively large units, such as paragraphs or entire chapters. While this improves the reliability of individual measurements, it also limits the number of available samples. At the same time, textual data often exhibit high dimensionality due to feature-rich representations, such as frequent words, $n$-grams, or syntactic constructions, leading to a severe imbalance between the number of samples and features. This issue is especially prevalent in stylometric analysis, where subtle stylistic differences are captured through sparse, high-dimensional feature spaces (Stamatatos, 2009). Given these constraints, clustering methods that rely on covariance estimation or strong parametric assumptions struggle to produce stable and interpretable results.
@@
To address these limitations, we introduce a novel entropy-driven soft clustering framework that identifies formulaic structures based on the log probability ($\log p$) of data points within a given distribution. The choice of $-\log p$ as our core metric stems from fundamental principles in information theory: the information content of an observation $x$, given a probability distribution $p(x)$, is defined as $-\log p(x)$, known as its self-information. This quantity directly captures the degree of predictability of $x$—highly predictable elements (i.e., those frequently occurring in structured, formulaic segments) have higher probabilities and thus lower self-information values, whereas less structured, unpredictable elements exhibit lower probabilities and correspondingly higher self-information values.
@@
Our framework integrates soft clustering weights (Peters et al., 2013), which adaptively re-weight samples based on their self-information contributions, mitigating the influence of sparse or noisy data points. This probabilistic approach ensures robust partitioning of text segments by capturing both localized structural patterns and broader stylistic trends. Unlike traditional clustering methods that assume specific data distributions – such as Gaussian mixture models – our information-based approach remains effective across a wide range of feature distributions, given an adequate choice of the score function, including multimodal and categorical datasets. Furthermore, by avoiding direct covariance inversion, our method enhances numerical stability and reduces susceptibility to the curse of dimensionality. Through this paradigm, we offer a flexible, data-driven solution for identifying formulaic structures in textual corpora while maintaining robustness in challenging high-dimensional settings.
The paper is structured as follows: In §2, we introduce the mathematical foundations of the information-based soft clustering framework, defining the auto-entropy distribution and formulating an optimization scheme for distinguishing formulaic from non-formulaic textual components in §3, we benchmark the method against existing clustering techniques on synthetic datasets, assessing its ability to classify structured and unstructured elements under controlled conditions, in §4 we apply the framework to the analysis of the biblical corpus, evaluating its capacity to differentiate hypothesized literary strata based on stylistic and formulaic properties, and in §5 we conclude.
@@ 2 Soft Clustering Algorithm
The following sections establish the mathematical foundations of the proposed information-based soft clustering framework and formulate an optimization scheme for distinguishing formulaic from non-formulaic text layers. In this framework, “soft clustering” refers to the relaxation techniques employed during optimization to avoid indeterminate solutions (e.g., Royer, 2017). This approach follows established clustering methodologies, where constraints are relaxed to transform discrete, hard assignment problems into continuous optimization problems, making them more tractable. An example of this is semidefinite relaxation in graph clustering, which reformulates discrete clustering problems as continuous ones, enabling more efficient optimization. By utilizing such relaxation techniques, our framework can more effectively identify structured patterns in text while maintaining flexibility in handling ambiguity in cluster assignments.
@@ 2.1 Self-Information as a Measure for Identifying Formulaic Structures
which quantifies how predictable or surprising an observation is. More predictable elements (those that occur frequently in structured, formulaic segments) have higher probabilities and thus lower $-\log p(x)$ values, while less predictable elements exhibit lower probabilities and higher $-\log p(x)$ values.
In the context of textual analysis, formulaic structures, by their nature, are repetitive and highly predictable, leading to lower $-\log p(x)$ values for formulaic segments. Our method directly measures the degree of structural predictability of each text segment, providing a granular, data-driven approach to distinguishing formulaic from non-formulaic patterns. Unlike existing information-based clustering methods, which rely on global variability measures, our approach evaluates the local predictability of individual elements, ensuring sensitivity to complex and varied linguistic structures.

### v2

# An Unsupervised Information-Theoretic Approach to Identifying Formulaic Clusters in Textual Data

@@ Abstract
Texts, whether literary or historical, exhibit structural and stylistic patterns shaped by their purpose, authorship, and cultural context. Formulaic texts, which are characterized by repetition and constrained expression, tend to differ in their information content (as defined by Shannon) compared to more dynamic compositions. Identifying such patterns in historical documents, particularly multi-author texts like the Hebrew Bible, provides insights into their origins, purpose, and transmission. This study aims to identify formulaic clusters: sections exhibiting systematic repetition and structural constraints, by analyzing recurring phrases, syntactic structures, and stylistic markers. However, distinguishing formulaic from non-formulaic elements in an unsupervised manner presents a computational challenge, especially in high-dimensional and sample-poor data sets where patterns must be inferred without predefined labels.
To address this, we develop an information-theoretic algorithm leveraging weighted self-information distributions to detect structured patterns in text. Our approach directly models variations in sample-wise self-information to identify formulaicity. By extending classical discrete self-information measures with a continuous formulation based on differential self-information in multivariate Gaussian distributions, our method remains applicable across different types of textual representations, including neural embeddings under Gaussian priors.
@@
F. Author]gideon.yoffe@mail.huji.ac.il
Plain Language Summary
This work presents a method for identifying formulaic clusters—segments of text that rely heavily on repetition, structural regularity, or stylistic constraint. Such clusters can be found in a wide range of contexts, including religious texts, technical manuals, oral traditions, and institutional writing, where language is shaped by convention, genre, or the need for consistency. Detecting these patterns can offer insight into how texts were composed, transmitted, and shaped by cultural forces over time.
Our approach is grounded in information theory and focuses on the concept of self-information, which measures how predictable or unpredictable a text unit is with respect to the corpus with which it is associated. Formulaic segments tend to be more predictable and, therefore, carry higher self-information. By measuring the distribution of self-information across a corpus, we identify areas where language use is more constrained, indicating a higher degree of formulaicity.
Benchmarking against standard clustering methods on synthetic and real data confirms the method’s robustness, particularly in high-dimensional, low-sample contexts. It outperforms conventional models in isolating internally consistent structures, without requiring strong distributional assumptions.
We apply our method to three books of the Hebrew Bible: Genesis, Exodus, and Leviticus. In Genesis, we focus on the distinction between genealogical lists and surrounding narrative material. The genealogical sections are known for their rigid, repetitive structure, and our method effectively isolates them as highly formulaic. In Exodus, we examine the division between Priestly (P) and non-Priestly texts, with the Priestly material comprising legal and cultic instructions. The method reliably separates these layers in line with traditional source-critical scholarship. In Leviticus, we test whether the method can distinguish between the Priestly source and the Holiness code (H), two stylistically distinct strata. The results show that subtle differences in recurring expressions and structural regularity allow for a clear separation between them under appropriate conditions.
This work offers a flexible, information-theoretic framework for analyzing the internal structure of texts. It is well suited to the study of layered, edited, or composite documents, where patterns of repetition and constraint signal deeper organizational principles. The method is broadly applicable to literary, historical, legal, or technical corpora, wherever formulaic language plays a role in shaping textual form.
@@ 1 Introduction
An intriguing challenge in textual analysis is distinguishing formulaic layers—those characterized by recurring patterns, motifs, or stylistic conventions—from non-formulaic ones. In this context, a formulaic element is defined as a word or a sequence thereof, syntactic or grammatical structures, or thematic patterns that recur systematically within a corpus, either due to conventionalized linguistic usage, genre-specific constraints, or cultural transmission mechanisms. Formulaic expressions can function as structural scaffolds within texts, shaping their composition and reflecting underlying conventions of specific literary traditions, genres, or modes of authorship (read2008measurement, 49, 40, 65, e.g.,). Beyond their structural significance, analyzing formulaic patterns provides insights into the dynamic interplay between creativity and constraint in language, illuminating the socio-cultural and cognitive forces that have influenced textual transmission and stylistic evolution (magoun1953oral, 38, 30, e.g.,).
@@
The importance of formulaicity is particularly pronounced in historical texts, where composition layers often reflect complex authorship and redaction processes. Many historical documents, created and transmitted over centuries, are products of collective memory and institutional influence. As such, these texts often contain formulaic structures that encode cultural norms, narrative conventions, and theological or political ideologies (knohl2007sanctuary, 33, 45, e.g.,). Understanding these layers can reveal how texts were constructed, transmitted, and adapted over time (gitay1980tradition, 21, 44, 11, e.g.,).
@@
Among historical texts, the Hebrew Bible provides a compelling case study. As a composite text shaped by centuries of authorship, redaction, and transmission, the Hebrew Bible embodies a rich interplay of formulaic and non-formulaic elements. Its stylistic diversity reflects contributions from different schools of authorship, genres, and literary motifs, ranging from tightly structured legal codes to more fluid poetic and narrative traditions warters1976formula (63, 55, 59). Detection and analysis of formulaic patterns within the Hebrew Bible offer a pathway to better understanding its compositional history, shedding light on how diverse voices and traditions coalesced into the canonical form.
@@
Understanding the distinction between formulaic and non-formulaic layers requires rigorous methodologies to quantify stylistic and structural patterns across literary feature sets. One promising metric for addressing these challenges is entropy, a concept from information theory that measures the unpredictability or randomness of a dataset. Initially introduced by Rudolf Clausius in 1865 as a measure of energy dispersal in a system, and subsequently formalized statistically by shannon1948mathematical (54), entropy has found applications in diverse fields, including thermodynamics, cryptography, and data compression. In the context of textual analysis, entropy provides a framework for quantifying the variability and regularity inherent in linguistic structures. Formally, given a discrete probability distribution $P=\{p_{1},p_{2},\dots,p_{n}\}$ over $n$ possible elements, Shannon entropy is defined as:
@@
where $p_{i}$ represents the probability of the $i$-th element. Higher entropy values indicate greater variability, while lower entropy values correspond to constrained, repetitive structures, making entropy a powerful tool for distinguishing formulaic clusters from more unpredictable distributions.
The relevance of entropy to textual analysis lies in its ability to capture the degree of predictability of literary or linguistic features within a sequence, such as a sentence, paragraph, or chapter. Texts with high formulaicity, such as legal codes or liturgical compositions, are likely to exhibit low entropy, such that text units within them resemble one another, for example, in vocabulary or grammatical structure. Conversely, more creative or extemporaneous texts often exhibit higher entropy, reflecting their greater unpredictability. Information-based metrics, therefore, serve as a natural tool for distinguishing between formulaic and non-formulaic elements within texts (church1990word, 9, e.g.,).
@@
where $w_{i}$ represents the $i$-th word in the sequence, and $P(w_{i}\mid w_{i-1},w_{i-2},\dots)$ denotes the conditional probability of $w_{i}$ given its preceding context. Perplexity has proven instrumental in analyzing textual corpora over various linguistic angles (klakow2002testing, 31, 26, 19, e.g.,). In the context of formulaic analysis, perplexity can reveal how tightly a text adheres to specific linguistic conventions (kurzynski2023stylometry, 35, e.g.,).
@@
For instance, Gaussian Mixture Models (GMMs) vlassis2002greedy (62) assume that the dataset can be represented as a weighted sum of Gaussian components, enabling efficient clustering through iterative optimization of parameters such as means, covariance matrices, and mixing weights. Similarly, Cross-Entropy Clustering (CEC) tabor2014cross (60) explicitly incorporates entropy as a criterion to partition data, refining clusters based on their statistical coherence. These methods, while effective in many contexts, rely on assumptions about the underlying data distribution, which may not always align with the structural and stylistic non-Gaussian variability inherent in textual corpora.
Recent advances in representation learning, including the Contrastive Language–Image Pretraining model (CLIP; radford2021clip (48)) and Bidirectional Encoder Representations from Transformers (BERT; devlin2018bert (14)), have introduced powerful mechanisms for unsupervised clustering and alignment in high-dimensional semantic spaces. These models often optimize mutual information across paired or contextualized inputs, effectively capturing latent structural relationships without requiring explicit probabilistic assumptions. Building on these architectures, recent work has proposed information-theoretic evaluation metrics computed over embedding spaces to quantify diversity and structural complexity. One such class of methods includes entropy measures based on Rényi Kernel Entropy (RKE), which estimate entropy non-parametrically using similarity kernels to capture the dispersion of representations in high-dimensional spaces principe2010information (46, 20). These approaches are increasingly used to assess representational diversity and clustering compactness in generative models and learned embeddings. Many of these entropy-based methods derive from or build upon neural embedding models such as variational autoencoders (VAEs), which assume that latent features follow a Gaussian distribution in a continuous space dilokthanakul2016deep (15, 66, 36). VAEs excel in capturing the underlying structure of data by encoding inputs into compact latent representations governed by a Gaussian prior casale2018gaussian (8), which enables smooth interpolations and the generation of new data points resembling the original distribution.
@@
While these approaches provide useful frameworks for clustering and representation learning, their reliance on Gaussian assumptions and metrics that require the estimation of covariance presents notable challenges. In textual data, where features often exhibit categorical, skewed, or multimodal distributions, the Gaussian prior may not accurately capture structural variability. Moreover, many of these methods are highly sensitive to the sample-to-dimension ratio altonji1996small (2, 67). When the sample size is small relative to the feature space, covariance estimation becomes unstable, leading to poorly conditioned matrices or singularities (ashurbekova2021optimal, 4, e.g.,). Regularization techniques, such as adding small positive constants to the diagonal of covariance matrices, can partially mitigate these issues but introduce biases that require careful tuning bickel2008regularized (5). This bias arises because regularization alters the estimated eigenvalues of the covariance matrix, effectively shrinking them toward a predetermined value, which can distort the true structure of the data and influence clustering or classification outcomes.
Crucially, while these techniques are effective in high-resource, general-purpose NLP settings, their applicability to low-resource or singular historical corpora remains limited. In domains such as ancient languages, where pretrained embeddings are unreliable or unavailable, and where interpretability is essential for source-critical and literary-historical inquiry, alternative approaches are required.
@@
These challenges are particularly pronounced in textual analysis, where statistical significance often necessitates dividing texts into relatively large units, such as paragraphs or entire chapters. While this improves the reliability of individual measurements, it also limits the number of available samples. At the same time, textual data often exhibit high dimensionality due to feature-rich representations, such as frequent words, $n$-grams, or syntactic constructions, leading to a severe imbalance between the number of samples and features. This issue is especially prevalent in stylometric analysis, where subtle stylistic differences are captured through sparse, high-dimensional feature spaces Stamatatos2009 (58). Given these constraints, clustering methods that rely on covariance estimation or strong parametric assumptions struggle to produce stable and interpretable results.
@@
To address these limitations, we introduce an entropy-driven soft clustering framework that identifies formulaic structures based on the log probability ($\log p$) of data points within a given distribution. The choice of $\log p$ as our core metric stems from fundamental principles in information theory: the information content of an observation $x$, given a probability distribution $p(x)$, is defined as $\log p(x)$, known as its self-information. This quantity directly captures the degree of predictability of xx—highly predictable elements (i.e., those frequently occurring in structured, formulaic segments) have higher probabilities and thus lower self-information values, whereas less structured, unpredictable elements exhibit lower probabilities and correspondingly higher self-information values.
@@
Our framework integrates soft clustering weights peters2013soft (42), which adaptively re-weight samples based on their self-information contributions, mitigating the influence of sparse or noisy data points. This probabilistic approach ensures robust partitioning of text segments by capturing both localized structural patterns and broader stylistic trends. Unlike traditional clustering methods that assume specific data distributions, such as Gaussian mixture models, our information-based approach remains effective across a wide range of feature distributions, given an adequate choice of the score function, including multimodal and categorical datasets. Furthermore, by avoiding direct covariance inversion, our method enhances numerical stability and reduces susceptibility to the curse of dimensionality. Through this paradigm, we offer a flexible, data-driven solution for identifying formulaic structures in textual corpora while maintaining robustness in challenging high-dimensional settings.
To accommodate different types of textual feature representations, we consider two complementary formulations of the method: one for discrete data and one for continuous data. The discrete formulation operates directly on binary or count-based features, avoiding any form of covariance estimation. In contrast, the continuous formulation, used for real-valued vector representations, such as embeddings, does involve estimating a soft covariance matrix. While this introduces some of the challenges seen in classical approaches, our method confines the estimation to weighted subsets of the data associated with candidate clusters exhibiting internal coherence. This restriction improves numerical stability by avoiding cross-cluster covariance terms, which are often the most ill-conditioned. Moreover, in both formulations, the objective function remains grounded in self-information, prioritizing structural predictability over explicit density modeling. This unifying information-theoretic perspective allows the method to remain robust across data modalities while still benefiting from statistical sensitivity to underlying stylistic regularities.
The paper is structured as follows: In §2, we introduce the mathematical foundations of the information-based soft clustering framework, defining the self-information distribution and formulating an optimization scheme for distinguishing formulaic from non-formulaic textual components in §3, we benchmark the method against existing clustering techniques on synthetic datasets, assessing its ability to classify structured and unstructured elements under controlled conditions, in §4 we apply the framework to the analysis of the biblical corpus, evaluating its capacity to differentiate hypothesized literary strata based on stylistic and formulaic properties, and in §5 we conclude.
@@ 2 Soft Clustering Algorithm
The following sections establish the mathematical foundations of the proposed information-based soft clustering framework and formulate an optimization scheme for distinguishing formulaic from non-formulaic text layers. In this framework, ”soft clustering” refers to the relaxation techniques employed during optimization to avoid indeterminate solutions (royer2017adaptive, 52, e.g.,). This approach follows established clustering methodologies, where constraints are relaxed to transform discrete, hard assignment problems into continuous optimization problems, making them more tractable. An example of this is semidefinite relaxation in graph clustering, which reformulates discrete clustering problems as continuous ones, enabling more efficient optimization. By utilizing such relaxation techniques, our framework can more effectively identify structured patterns in text while maintaining flexibility in handling ambiguity in cluster assignments.
@@ 2.1 Self-Information as a Measure for Identifying Formulaic Structures
which quantifies how predictable or surprising an observation is. More predictable elements (those that occur frequently in structured, formulaic segments) have higher probabilities and thus lower $\log p(x)$ values, while less predictable elements exhibit lower probabilities and higher $\log p(x)$ values.
In the context of textual analysis, formulaic structures, by their nature, are repetitive and highly predictable, leading to higher $\log p(x)$ values for formulaic segments. Our method directly measures the degree of structural predictability of each text segment, providing a granular, data-driven approach to distinguishing formulaic from non-formulaic patterns. Unlike existing information-based clustering methods, which rely on global variability measures, our approach evaluates the local predictability of individual elements, ensuring sensitivity to complex and varied linguistic structures.
Layman explanation:
Self-information provides a measure of how surprising or expected a given unit of text is, based on how likely it is to appear. Frequent and structurally constrained expressions are more predictable and therefore contribute higher self-information values. By measuring this quantity locally for each segment, our method identifies areas of the text that exhibit heightened regularity or constraint, which are often associated with formulaic language.